The city of Raleigh has 11,400 registered voters. There are two candidates for city council in an upcoming election: Brown and Feliz. The day before the election, a telephone poll of 200 randomly selected registered voters was conducted. 84 said they'd vote for Brown, 108 said they'd vote for Feliz, and 8 were undecided.

Give the sample statistic for the proportion of voters surveyed who said they'd vote for Brown. Round your answer to four decimal places. to find answer

  

This sample statistic suggests that we might expect how many of the 11,400 registered voters to vote for Brown?
Answer #1

a)

sample statistic for the proportion of voters surveyed who said they'd vote for Brown.
pcap = 84/200 = 0.4200

b)
expected number of votes = 11400 * 0.42 = 4788